jack said , five times my age in two years is 100, how old is jack now?
i need help on writing an equation to get the answer 20(: thanks so much!

Answer by Maths68(1474)

Let
Present age of Jack = x
In two years he will be = x+2
Five time in two years means = 5(x+2)
So
5(x+2)=100
5x+10=100
5x=100-10
5x=90
5x/5=90/5
x=18
Jack is 18 years old Now after two years he will be 20 and 5 times of 20 is 100